### v1

@@ 2.2 Our Contribution
A tree cut-sparsifier of $G$ of quality $\alpha$ is also a tree flow-sparsifier of quality $O(\alpha\log n)$. This follows from the flow-cut gap theorem [leighton1999multicommodity, aumann1998log] (stated below), together with the fact that the flow-cut gap in a tree is $1$.
Theorem 2.10 (Flow-Cut Gap [aumann1998log]).
Let $Q$ be a multi-commodity flow problem on $G=(V,E)$, with demand matrix $D_{Q}$. Define $\theta_{G}(Q)\ :=\ \sup\{\theta:\text{$G$ $\theta$-respects $Q$}\}.$ Then, for a universal constant $c>0$,
Using Theorem 2.10, we get that a tree cut-sparsifier $T$ of $G$ of quality $\alpha$ satisfies, for all flow problems $Q$,
$\operatorname{cong}_{T}(Q)\leq\operatorname{cong}_{G}(Q)\leq\frac{\log n}{c\cdot\theta_{G}(Q)}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu1\mkern-1.5mu)}}{{\leq}}\frac{\alpha\log n}{c\cdot\theta_{T}(Q)}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{=}}\frac{\alpha\log n}{c}\operatorname{cong}_{T}(Q),$
where Inequality (1) holds since $T$ is a cut-sparsifier of quality $\alpha$, and Equality (2) holds since the flow-cut gap of a tree is 1. Therefore, $T$ is a flow-sparsifier of $G$ of quality $O(\alpha\log n)$.
@@
The flow-cut gap for planar graphs is $O(\sqrt{\log n})$ [rao1999small]. Therefore, for planar graphs, our construction achieves tree flow-sparsifier (and congestion-approximator) of quality $O(\log^{2.5}n\log\log n)$.
@@ 2.3 Certifying a Tree Cut-Sparsifier Versus Certifying a Tree Flow-Sparsifier
Our construction follows a recursive partitioning algorithm which takes as input a cluster $S\subseteq V$ and partitions it to sub-clusters $S_{1},\ldots,S_{t}$. Applying this algorithm recursively gives a laminar decomposition of the set $V$, which naturally corresponds to a tree $T$. The root of $T$ corresponds to the set $V$ while the leaves correspond to individual vertices $v\in V$. As is standard in tree flow-sparsifier algorithms, the weight of an edge $(S,P(S))$ in $T$ between the node corresponding to the cluster $S$ and its parent $P(S)$ is $\operatorname{cap}_{G}(S,V\setminus S)$.
@@
Consider first the task of certifying that such a tree $T$ is a flow-sparsifier. We can do this by showing that the congestion of routing a demand matrix $Q$ in $T$ is comparable to the congestion of routing $Q$ in $G$. Note that by choosing the capacities of the tree edges as we did, the inequality $\operatorname{cong}_{T}(Q)\leq\operatorname{cong}_{G}(Q)$ is immediate: we route the multi-commodity flow problem $Q$ in $T$ trivially along the unique path between each pair of vertices. If the flow problem can be routed in $G$ with congestion $c$, then by Fact 2.7 it must be that, for every sub-cluster $S\in V_{T}$, the cut $(S,V\setminus S)$ $c$-respects $Q$: $\frac{\operatorname{cap}(S,V\setminus S)}{\operatorname{dem}_{Q}(S,V\setminus S)}\geq c$. The demand across each cut $(S,V\setminus S)$ corresponds exactly to the total amount of flow routed along the edge $(S,P(S))$ in the tree, so this ensures that the flow problem can be routed in the tree with congestion $c$. Therefore, when building a tree flow-sparsifier, our goal is to choose the tree such that the reverse inequality $\operatorname{cong}_{G}(Q)\leq\alpha\cdot\operatorname{cong}_{T}(Q)$ also holds. This means that given a demand matrix $Q$ which can be routed in $T$ we have to show how to route it in $G$ with similar congestion. $Q$ being routable in $T$ means that all cuts $(S,V\setminus S)$, for $S\in V_{T}$, respect $Q$. Therefore, the intuition is that we look to include in the tree the sub-clusters which best capture the cut structure of the graph, in the sense that any demand matrix $Q$ respected by them can be routed in $G$.
@@
In the flow-sparsifier case, given a demand matrix $Q$ which is routable in $T$, [harrelson2003polynomial, bienkowski2003practical, racke2014computing] used the tree structure to describe the routing of $Q$ in the graph $G$. They did this by routing $Q$ which originates at all the vertices of $V$ (corresponding to leaves in $T$), toward the boundaries of sub-clusters $S$ as they go up along the tree. [^2] This is made formal using the subdivision graph (see Definition 3.1), where the demands will reside on the split nodes of the boundary edges. For the overview in this section, however, we omit these details for simplicity. That is, when a sub-cluster $S\in T$ is partitioned into $\{S_{1},\ldots,S_{t}\}$, they route the mass in $G$, from the boundary edges of $\{S_{i}\}_{i=1}^{t}$ (where they arrived in the previous iteration) to the boundary edges of $S$. We emphasize that this routing takes place in $G$, while the structure of $T$ is used only as a “blueprint” to direct the intermediate flows. They then “pair-up” source and target demands of vertex pairs that are both in $S$ before moving further upwards in the tree.
@@
Turning now back to cut-sparsifiers, in order to certify that $T$ is a tree cut-sparsifier, we no longer need to route the demand matrix $Q$ (which can be routed in $T$) in $G$, but rather only show that $G$ respects $Q$ (see Claim 5.2 which essentially says that $T$ is a tree cut-sparsifier of quality $\alpha$ if it $1/\alpha$-respects every demand matrix $Q$ that is $1$-respected by $T$). To show that demand matrix $Q$ is respected by $G$ we develop a technique of moving demands rather than flow. For this we define the notion of a valid demand state (see Definition 3.13).
@@
We can then update the demand state by moving mass around, thus reasoning about intermediate steps in the routing scheme. That is, we again use the structure of $T$ to direct the movement of demands. As we climb up the sparsifier $T$, we move the demands of the demand state $P$ from internal edges to the boundary edges of sub-clusters $S\in V_{T}$. However, unlike in the flow-sparsifier setting, these movements do not necessarily correspond to an actual multi-commodity flow that is routable in the graph. In fact, we do not “pay” for the congestion of these routings, as in the flow-sparsifier analysis, but rather according to how much the graph respects them (see Definition 2.6). This charging scheme is detailed in Section 5.
@@ 2.4 Dealing With Bottleneck Cuts
However, as discussed in [bienkowski2003practical, harrelson2003polynomial, RackeS14, racke2014computing], there may be sub-clusters $S\subseteq V$, for which we cannot find a partition $\{S_{i}\}_{i=1}^{t}$ such that $F\cup B$ is flow-linked in $G[S]$. A simple counter-example is that it might be the case that $B$ itself is not flow-linked in $G[S]$, or maybe even disconnected there. In this case, no partition of $G[S]$ will work. We use the name bottleneck cut to refer to a cut which certifies that the boundary edges are not flow-linked. This name was introduced by [RackeS14], though their bottleneck cuts only had to satisfy a weaker requirement. A similar (yet not equivalent) notion appeared earlier in [bienkowski2003practical, harrelson2003polynomial]. The algorithms of [bienkowski2003practical, harrelson2003polynomial, RackeS14] had to make sure that the clusters have no bottleneck cuts. They achieved this by cutting clusters that have bottleneck cuts. We call this step the refinement phase [RackeS14]. Previous work implement the refinement phase in different ways, either by ensuring a “precondition” on the sub-clusters they recur into [bienkowski2003practical], or encountering these cuts and handling them “on the fly” [harrelson2003polynomial, RackeS14]. In either case dealing with these cuts slowed down these early algorithms substantially.
@@
Since ensuring that there is no bottleneck cut in each sub-cluster before recurring into it may be expensive (in terms of running time), [racke2014computing] settled for something slightly weaker: They first ignore the boundary edges of $S$ and find a partition of $S$ into $\{S_{1},\ldots,S_{t}\}$ such that the inter-cluster edges $F$ are flow-linked. Then they find an approximate min-cut $Y$ for the (single commodity) flow problem which routes flow from $F$ to $B$. By adding to $T$ a new sub-cluster whose boundary is $Y$, [racke2014computing] are able to bound the amount of mass that has to cross the cut $Y$, since now the demand matrix $Q$ is respected by $Y$ in $G$.
@@
Adding $Y$ as a boundary of a sub-cluster in $T$ is enough to show how to move the demands to the boundary edges $B$, such that the load on each boundary edge is multiplied by $1+\uptau$ units ($\uptau$ is a parameter $0<\uptau\leq 1$). However, the distribution of demands on the boundary edges is not uniform. Hence, demands whose source and target are in the sub-cluster do not cancel-out. Consequently, the load accumulates when going up along the tree, multiplied by a factor of at most $1+\uptau$ in each iteration. Choosing $\uptau=\frac{1}{\log n}$, such that the load is multiplied by a factor of $(1+\frac{1}{\log n})$ along the $O(\log n)$ levels of the tree, ensures that the load is constant. This method, however, incurs a degradation of $O(\frac{1}{\uptau})$ per level in the quality of the tree flow-sparsifier, leading to an overall loss of $O(\frac{\log n}{\uptau})=O(\log^{2}n)$. [^3] Choosing $\uptau=\frac{1}{\log n}$ is optimal, up to a multiplicative constant. Let $k=\Theta(\log n)$ be the depth of the tree. We omit the proof of the following claim: For any sequence of $\uptau_{1},\ldots,\uptau_{k}\leq 1$, it holds that $\sum_{i=1}^{k}{\frac{1}{\uptau_{i}}\prod_{j=i+1}^{k}{(1+\uptau_{j})}}=\Omega(k^{2})$. The additional $O(\log^{2}n)$ factor in their quality comes from the usage of [khandekar2009graph]’s cut-matching game for flow (so in total the quality is $O(\log^{4}n)$).
@@
In our first step below (Section 2.5) we deal with bottleneck cuts similarly to [racke2014computing], but using primitives from the expander decomposition algorithm in [ADK25], which are the cut-analogue of [khandekar2009graph]’s cut-matching game. This allows us to get Theorem 2.9. Afterwards, in Section 2.6 we describe a better way to deal with bottleneck cuts, by implementing a version of the refinement phase in near-linear time. This allows us to improve the quality of the resulting tree cut-sparsifier and get Theorem 2.8.
@@ 2.5 A Tree Cut-Sparsifier of Quality $O(\log^{3}n)$
We now adapt [racke2014computing] to the cut-sparsifier setting. As outlined in Section 2.3, our goal here is not to route the demand matrix $Q$ within $G$, but rather to ensure that $G$ respects $Q$. In this formulation, the analogue of a set $F$ of inter-cluster edges (within a cluster $S$) being flow-linked in [racke2014computing] is that $G[S]$ respects the all-to-all demand matrix between edges of $F$ (see Definition 2.2), or equivalently that $F$ expands in $G[S]$ (see Definition 3.5 below). Therefore, our first goal is to obtain, efficiently, such a set $F$. For this step we ignore the boundary edges of $S$ (which may be separated by bottleneck cuts). We use the following algorithmic tool:
@@
To obtain such an $F$ efficiently we use the “balanced cut or expander” primitive from [ADK25], which allows us to certify that the sub-cluster $G[S]$ respects the all-to-all flow problem on a subset $F\subseteq E$, or find a balanced cut which refutes this. We start with $F_{1}=E(G[S])$ and iterate on $i=1,\ldots$:
@@
This step is the cut-analogue of [racke2014computing, Lemma 3.1], but powered by [ADK25]’s tool so that the certification we obtain when $F$ is good is expansion (according to Definition 3.5) rather than being flow-linked.
@@
Once $F$ is fixed, we connect $F$ to the boundary edges similarly to [racke2014computing]. We compute a cut $Y$ (via an auxiliary approximate flow problem) that separates $F$ from $B$. For this purpose, we use the fair-cuts algorithm of [LNPSsoda23], which serves as an efficient approximate max-flow solver and provides slightly better congestion (see Lemma 4.2) compared to [racke2014computing, Lemma 3.2], as well as simplifying the argument.
The usage of [ADK25] rather than a procedure inspired by the cut-matching game of [khandekar2009graph] in [racke2014computing] allows us to improve the quality of the cut-sparsifier by a $\log n$ factor. Specifically, we still pay $O(\frac{\log n}{\uptau})=O(\log^{2}n)$ like in Section 2.4, but the additional factor from the cut-matching component is now $O(\log n)$ instead of $O(\log^{2}n)$. This gives a tree cut-sparsifier of quality of $O(\log^{3}n)$, and establishes Theorem 2.9. In particular, as discussed after Theorem 2.9, a tree cut-sparsifier of quality $\alpha$ implies a tree flow-sparsifier of quality $O(\alpha\log n)$. Therefore, this tree is also a tree flow-sparsifier of quality $O(\log^{4}n)$, matching the quality of the flow-sparsifier of [racke2014computing].
The construction of the tree cut-sparsifier is detailed in Section 4. Importantly, we use the modular building blocks of the balanced sparse cut procedure [ADK25] and the fair-cuts algorithm [LNPSsoda23] as a black box. This significantly simplifies the construction of the tree cut-sparsifier, when compared to, e.g., [racke2014computing], as it abstracts away many details of the underlying algorithmic tools.

### v2

@@ 2.2 Our Contribution
Remark 2.10 (Hierarchical trees).
The trees produced in Theorems 2.8 and 2.9 are hierarchical trees: each node of $T$ corresponds to a subset of vertices of $G$, and these subsets form a laminar family [^2] That is, the subsets corresponding to the children of a node in $T$ are pair-wise disjoint and contained in the subset corresponding to that node. in which the root corresponds to $V$ and the leaves correspond to all singleton subsets $\{v\}_{v\in V}$. Each edge $e\in E_{T}$ connects a node, corresponding to a cluster $S_{e}\subseteq V$, with its parent node, corresponding to a larger cluster $S_{e}\subseteq P_{T}(S_{e})\subseteq V$. This edge has a capacity equal to the capacity of the cut $(S_{e},V\setminus S_{e})$ in $G$, i.e., $\operatorname{cap}_{T}(e)=\operatorname{cap}_{G}(S_{e},V\setminus S_{e})$.
The following claim follows from the flow-cut gap theorem [19, 4], together with the fact that the flow-cut gap in a tree is $1$ [24]. We defer the (standard) proof to appendix A.
Claim 2.11.
The following relations hold for a (hierarchical) tree $T$ of $G$:
  • If $T$ is a cut-sparsifier of $G$ of quality $\alpha$, then $T$ is also a flow-sparsifier of quality $O(\alpha\log n)$.
  • If $T$ is a flow-sparsifier of $G$ of quality $\alpha$, then $T$ is also a cut-sparsifier of quality $\alpha$.
@@
The flow-cut gap for planar graphs is $O(\sqrt{\log n})$ [31]. Therefore, for planar graphs, our construction achieves tree flow-sparsifier (and congestion-approximator) of quality $O(\log^{2.5}n\log\log n)$.
@@ 2.3 Certifying a Tree Cut-Sparsifier Versus Certifying a Tree Flow-Sparsifier
Our construction follows a recursive partitioning algorithm which takes as input a cluster $S\subseteq V$ and partitions it to sub-clusters $S_{1},\ldots,S_{t}$. Applying this algorithm recursively gives a laminar decomposition of the set $V$, which naturally corresponds to a hierarchical tree $T$ (see Remark 2.10). The root of $T$ corresponds to the set $V$ while the leaves correspond to individual vertices $v\in V$. The weight of an edge $(S,P_{T}(S))$ in $T$, between the node corresponding to the cluster $S$ and its parent $P_{T}(S)$ is $\operatorname{cap}_{G}(S,V\setminus S)$.
@@
Consider first the task of certifying that such a tree $T$ is a flow-sparsifier. We can do this by showing that the congestion of routing a demand matrix $Q$ in $T$ is comparable to the congestion of routing $Q$ in $G$. Note that by choosing the capacities of the tree edges as we did, the inequality $\operatorname{cong}_{T}(Q)\leq\operatorname{cong}_{G}(Q)$ is immediate: we route the demand matrix $Q$ in $T$ trivially along the unique path between each pair of vertices. If the flow problem can be routed in $G$ with congestion $c$, then by Fact 2.7 it must be that, for every sub-cluster $S\in V_{T}$, the cut $(S,V\setminus S)$ $\frac{1}{c}$-respects $Q$: $\frac{\operatorname{cap}(S,V\setminus S)}{\operatorname{dem}_{Q}(S,V\setminus S)}\geq\frac{1}{c}$. The demand across each cut $(S,V\setminus S)$ corresponds exactly to the total amount of flow routed along the edge $(S,P_{T}(S))$ in the tree, so the way the edge capacities in $T$ are assigned ensures that the flow problem can be routed in the tree with congestion $c$. Therefore, when building a tree flow-sparsifier, our goal is to choose the tree such that the reverse inequality $\operatorname{cong}_{G}(Q)\leq\alpha\cdot\operatorname{cong}_{T}(Q)$ also holds. This means that given demands $Q$ which can be routed in $T$ we have to show how to route $Q$ in $G$ with similar congestion. $Q$ being routable in $T$ means that only for every $S\in V_{T}$, the cut $(S,V\setminus S)$ in $G$, respects $Q$. Therefore, the intuition is that we look to include in the tree the sub-clusters which best capture the cut structure of the graph, in the sense that any demands respected by them can be routed in $G$.
@@
In the flow-sparsifier case, given demands $Q$ routable in $T$, [15, 7, 27] used the tree structure to describe the routing of $Q$ in the graph $G$. They did this by routing $Q$ which originates at all the vertices of $V$ (corresponding to leaves in $T$), toward the boundaries of sub-clusters $S$ as they go up along the tree. [^3] This is made formal using the subdivision graph (see Definition 3.1), where the demands reside on the split nodes of the boundary edges. For the overview in this section, however, we omit these details for simplicity. That is, when a sub-cluster $S\in T$ is partitioned into $\{S_{1},\ldots,S_{t}\}$, they route (in $G$) the mass from the boundary edges of $\{S_{i}\}_{i=1}^{t}$ (where they arrived in the previous iteration) to the boundary edges of $S$. We emphasize that this routing takes place in $G$, while the structure of $T$ is used only as a “blueprint” to direct the intermediate flows. Then, source and target demands of vertex pairs that are both in $S$, are “paired up” and canceled before moving further upwards in the tree.
@@
Turning now back to cut-sparsifiers, in order to certify that $T$ is a tree cut-sparsifier, we no longer need to route the demands $Q$ (which can be routed in $T$) in $G$, but rather only show that $G$ respects $Q$ (see Claim 5.2 which essentially says that $T$ is a tree cut-sparsifier of quality $\alpha$ if it $1/\alpha$-respects every demands $Q$ that are $1$-respected by $T$). To show that demands $Q$ are respected by $G$ we develop a technique of moving demands rather than flow. For this we define the notion of a valid demand state (see Definition 3.12).
@@
We use the structure of $T$ to direct the movement of demands. This corresponds to updating the demand state. Specifically, as we climb up the sparsifier $T$, we move the demands of the demand state from internal edges to the boundary edges of sub-clusters $S\in V_{T}$. However, unlike in the flow-sparsifier setting, these movements do not necessarily correspond to an actual multi-commodity flow that is routable in the graph. In fact, we do not “pay” for the congestion of these routings, as in the flow-sparsifier analysis, but rather according to how much the graph respects them (see Definition 2.6). This charging scheme is detailed in Section 5.
@@ 2.4 Dealing With Bottleneck Cuts
However, as discussed in [7, 15, 28, 27], there may be sub-clusters $S\subseteq V$, for which we cannot find a partition $\{S_{i}\}_{i=1}^{t}$ such that $F\cup B$ is flow-linked in $G[S]$. A simple counter-example is that it might be the case that $B$ itself is not flow-linked in $G[S]$, or maybe even disconnected there. In this case, no partition of $G[S]$ will work. We use the name bottleneck cut to refer to a cut which certifies that the boundary edges are not flow-linked. This name was introduced by [28], though their bottleneck cuts only had to satisfy a weaker requirement. A similar (yet not equivalent) notion appeared earlier in [7, 15]. The algorithms of [7, 15, 28] had to make sure that the clusters have no bottleneck cuts. They achieved this by cutting clusters that have bottleneck cuts. We call this step the refinement phase [28]. Previous work implement the refinement phase in different ways, either by ensuring a “precondition” on the sub-clusters they recur into [7], or encountering these cuts and handling them “on the fly” [15, 28]. In either case dealing with these cuts slowed down these early algorithms substantially.
@@
Since ensuring that there is no bottleneck cut in each sub-cluster before recurring into it may be expensive (in terms of running time), [27] settled for something slightly weaker: They first ignore the boundary edges of $S$ and find a partition of $S$ into $\{S_{1},\ldots,S_{t}\}$ such that the inter-cluster edges $F$ are flow-linked. Then they find an approximate min-cut $Y$ for the (single commodity) flow problem which routes flow from $F$ to $B$. By adding to $T$ a new sub-cluster whose boundary is $Y$, [27] are able to bound the amount of mass that has to cross the cut $Y$, since now the demand matrix $Q$ is respected by $Y$ in $G$.
@@
Adding $Y$ as a boundary of a sub-cluster in $T$ is enough to show how to move the demands to the boundary edges $B$, such that the load on each boundary edge is multiplied by $1+\uptau$ units ($\uptau$ is a parameter $0<\uptau\leq 1$). However, the distribution of demands on the boundary edges is not uniform. Hence, demands whose source and target are in the sub-cluster do not cancel-out. Consequently, the load accumulates when going up along the tree, multiplied by a factor of at most $1+\uptau$ in each iteration. Choosing $\uptau=\frac{1}{\log n}$, such that the load is multiplied by a factor of $(1+\frac{1}{\log n})$ along the $O(\log n)$ levels of the tree, ensures that the load is constant. This method, however, incurs a degradation of $O(\frac{1}{\uptau})$ per level in the quality of the tree flow-sparsifier, leading to an overall loss of $O(\frac{\log n}{\uptau})=O(\log^{2}n)$. [^4] Choosing $\uptau=\frac{1}{\log n}$ is optimal, up to a multiplicative constant. Let $k=\Theta(\log n)$ be the depth of the tree. We omit the proof of the following claim: For any sequence of $\uptau_{1},\ldots,\uptau_{k}\leq 1$, it holds that $\sum_{i=1}^{k}{\frac{1}{\uptau_{i}}\prod_{j=i+1}^{k}{(1+\uptau_{j})}}=\Omega(k^{2})$. The additional $O(\log^{2}n)$ factor in their quality comes from the usage of [18]’s cut-matching game for flow (so in total the quality is $O(\log^{4}n)$).
@@
In our first step below (Section 2.5) we deal with bottleneck cuts similarly to [27], but using primitives from the expander decomposition algorithm in [2], which are the cut-analogue of [18]’s cut-matching game. This allows us to get Theorem 2.9. Afterwards, in Section 2.6 we describe a better way to deal with bottleneck cuts, by implementing a version of the refinement phase in near-linear time. This allows us to improve the quality of the resulting tree cut-sparsifier and get Theorem 2.8.
@@ 2.5 A Tree Cut-Sparsifier of Quality $O(\log^{3}n)$
We now adapt [27] to the cut-sparsifier setting. As outlined in Section 2.3, our goal here is not to route the demand matrix $Q$ within $G$, but rather to ensure that $G$ respects $Q$. In this formulation, the analogue of a set $F$ of inter-cluster edges (within a cluster $S$) being flow-linked in [27] is that $G[S]$ respects the all-to-all demand matrix between edges of $F$ (see Definition 2.2), or equivalently that $F$ expands in $G[S]$ (see Definition 3.5 below). Therefore, our first goal is to obtain, efficiently, such a set $F$. For this step we ignore the boundary edges of $S$ (which may be separated by bottleneck cuts). We use the following algorithmic tool:
@@
To obtain such an $F$ efficiently we use the “balanced cut or expander” primitive from [2], which allows us to certify that the sub-cluster $G[S]$ respects the all-to-all flow problem on a subset $F\subseteq E$, or find a balanced cut which refutes this. We start with $F_{1}=E(G[S])$ and iterate on $i=1,\ldots$:
@@
This step is the cut-analogue of [27, Lemma 3.1], but powered by [2]’s tool so that the certification we obtain when $F$ is good is expansion (according to Definition 3.5) rather than being flow-linked.
@@
Once $F$ is fixed, we connect $F$ to the boundary edges similarly to [27]. We compute a cut $Y$ (via an auxiliary approximate maximum flow problem) that separates $F$ from $B$. For this purpose, we use the fair-cuts algorithm of [21], which serves as an efficient approximate max-flow solver and provides slightly better congestion (see Lemma 4.2) compared to [27, Lemma 3.2], as well as simplifying the argument.
The usage of [2] rather than a procedure inspired by the cut-matching game of [18] in [27] allows us to improve the quality of the cut-sparsifier by a $\log n$ factor. Specifically, we still pay $O(\frac{\log n}{\uptau})=O(\log^{2}n)$ like in Section 2.4, but the additional factor from the cut-matching component is now $O(\log n)$ instead of $O(\log^{2}n)$. This gives a tree cut-sparsifier of quality of $O(\log^{3}n)$, and establishes Theorem 2.9. In particular, as discussed after Theorem 2.9, a tree cut-sparsifier of quality $\alpha$ implies a tree flow-sparsifier of quality $O(\alpha\log n)$. Therefore, this tree is also a tree flow-sparsifier of quality $O(\log^{4}n)$, matching the quality of the flow-sparsifier of [27].
The construction of the tree cut-sparsifier is detailed in Section 4. Importantly, we use the modular building blocks of the balanced sparse cut procedure [2] and the fair-cuts algorithm [21] as a black box. This significantly simplifies the construction of the tree cut-sparsifier, when compared to, e.g., [27], as it abstracts away many details of the underlying algorithmic tools.